OpenStudy (anonymous):

cos 4x + cos 2x = 2 - 2 sin^2 2x - 2 sin^2 x verify the identity. I really dont understand this so if you can show steps to it too thatd be great:)

OpenStudy (ajprincess):

Take the identities on the left side. cos2x=1-2sin^2x cos4x=cos2(2x)=1-2sin^2(2x)

OpenStudy (ajprincess):

Getting it @beachgirl129?

OpenStudy (anonymous):

um I think like i see how you did those steps but once I get 1-2sin^2(2x) how will i get it to equal the right

OpenStudy (ajprincess):

cos4x+cos2x=1-2sin^2(2x)+1-2sin^2x

OpenStudy (anonymous):

ohhhkay thanks:)

OpenStudy (ajprincess):

welcome:)